OpenStudy (anonymous):

How could you eliminate the fractions before proceeding to solve this equation? five-eighths equals three-eighths times t plus one-eighth Add one-eighth to both sides of the equation. Subtract one-eighth from both sides of the equation. Divide both sides of the equation by 8. Multiply both sides of the equation by 8.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

5/8 = 3/8t + 1/8 (multiply the entire equation by the LCD, which is 8) 5 = 3t + 1 -- subtract 1 form both sides 5 - 1 = 3t 4 = 3t -- divide by 3 4/3 = t

OpenStudy (anonymous):

when you multiply the equation by the LCD, then the fractions turn to whole numbers

OpenStudy (anonymous):

ok so where multiplying by both sides

OpenStudy (anonymous):

yes...the entire equation

OpenStudy (anonymous):

ok